### v1

# Needle: A Generative-AI Powered Monte Carlo Method for Answering Complex Natural Language Queries on Multi-modal Data

Abstract.
Multi-modal data, such as image data sets, often miss the detailed descriptions that properly capture the rich information encoded in them. This makes answering complex natural language queries a major challenge in these domains. In particular, unlike the traditional nearest-neighbor search, where the tuples and the query are modeled as points in a data cube, the query and the tuples are of different natures, making the traditional query answering solutions not directly applicable for such settings. Existing literature addresses this challenge for image data through vector representations jointly trained on natural language and images. This technique, however, underperforms for complex queries due to various reasons.
This paper takes a step towards addressing this challenge by introducing a Generative-AI (GenAI) powered Monte Carlo method that utilizes foundation models to generate synthetic samples that capture the complexity of the natural language query and transform it to the same space of the multi-modal data. Following this method, we develop a system for image data retrieval and propose practical solutions that enable leveraging future advancements in GenAI and vector representations for improving our system’s performance. Our comprehensive experiments on various benchmark datasets verify that our system significantly outperforms state-of-the-art techniques.
Image Databases; Generative AI; Foundation Models;
1. Introduction
a Needle in a haystack: something that is impossible or extremely difficult to find, especially because the area you have to search is too large.”
Cambridge Dictionary
Figure 1. The query results for the queries: (a) [a banana] (row 1), (b) [an unripe banana] (row 2), and (c) [a banana gazing at its reflection in a mirror] (row 3), using Needle (first column) v.s. CLIP (second column).
Multi-modal data, including image datasets, pose new challenges for data management systems. Unlike traditional databases, where tuple attributes are explicitly specified, multi-modal data often miss proper descriptions that capture the rich information encoded in them. As a result, traditional query-answering approaches are not directly applicable in these settings. On the other hand, advanced personal devices such as cell phones have enabled the collection of large volumes of multi-modal data, even at the individual user level. Due to the wealth of information “hidden” in each multi-modal tuple, the users may compose complex natural language queries, describing the tuple(s) they are interested in. To further motivate this, let us consider the following running example.
Example 1.1.
A photo enthusiast has collected a large pool of images over the past several years on her private data repository. She wants to retrieve a specific photo she has in mind. She describes the picture as [A banana gazing at its reflection in a mirror].
Traditional nearest-neighbor (NN-) search algorithms are not directly applicable here since those consider the database tuples and the query as points in a data cube. In contrast, in Example 1.1, (i) each tuple is an image, and (ii) the query is in the form of an unstructured, complex natural language statement.
The state-of-the-art approach for addressing the need for answering natural language queries on image datasets is CLIP (Radford et al., 2021), a vector representation (aka embedding) jointly trained on several hundred million (image, text) pairs. Using CLIP as the embedder, one can transform the query and the dataset images to vector representation in the same space and apply NN-search to answer the query. This approach works satisfactorily for “simple” queries, such as simple object detection tasks. To see a concrete example, following with our running example (Example 1.1), we used a benchmark image dataset (LVIS (Gupta et al., 2019)) and issued the query [A banana]. As reflected in the top-right (row 1, col. 2) cell of Figure 1, the results were satisfactory, as most of the returned results include a banana (it also returned some irrelevant results, such as an image of a person wearing a banana costume). However, it failed for a moderately complex query, [An unripe banana], as none of the returned images was an unripe (green) banana – the middle-right (row 2, col. 2) cell of Figure 1. It further failed for the more complex queries, including the one in Example 1.1, and was not able to find the image of interest in its results. – the bottom-right (row 3, col. 2) cell of Figure 1. As we shall further explain in § 6, we observed a consistent behavior for this approach across a wide range of queries.
This motivates us to take a step toward filling this gap in this paper by proposing a Monte Carlo randomized algorithm. In particular, we empower our algorithm with generative AI (GenAI) to generate a set of synthetic multi-modal tuples that represent the complex natural language query. Then, leveraging a collection of embedders, we conduct a series of NN-search using the generated tuples and aggregate the results to generate the final query outputs. We use our Monte Carlo method to develop Needle, our system for complex natural query answering on image datasets.
Unlike existing approaches, Needle is capable of handling natural language queries with different levels of complexity. To demonstrate this, let us consider our running example, Example 1.1 once again. To compare Needle with CLIP, we issued the same set of simple, moderate, and complex queries. The results are provided in the first column of Figure 1. For the simple object-finding task, the top-left (row 1, col 1) cell in Figure 1, we observed a similar (but slightly better) performance since all returned images satisfied the query, i.e., they included a banana. It also performed reasonably well for the moderately complex query, as a large portion of the output images were green bananas – the middle-left (row 2, col 1) cell in Figure 1. But perhaps, more interestingly, it could successfully find the corresponding image described in the complex query of Example 1.1 – the bottom-left (row 3, col 1) cell in Figure 1. Interestingly, the target image was returned as the top-1 search result (the top-left image in the bottom-left cell), demonstrating the ability of our system to answer these types of queries. We observed a consistent behavior for our system in our experiments across a large number of queries (§ 6).
Summary of contributions
This paper introduces a new approach to using GenAI for complex natural-language query answering on multi-modal data. Accordingly, it proposes Needle, a system for image data retrieval. In summary, our contributions are the following:
  • We introduce a novel Monte Carlo method for answering complex natural-language queries on multi-modal data (§ 3). We leverage Foundation models to generate synthetic tuples representing natural language queries. We then use a collection of vector representations (aka embeddings) that allows us to apply traditional k-NN techniques to retrieve related tuples. Our Monte Carlo algorithm aggregates the output generated for each (synthetic tuple, embedding) pair to obtain the final result.
  • We use our Monte Carlo method to develop Needle, our interactive system image data retrieval (§ 4).
  • We propose a tiling technique, Smart Tiling, inspired by quadtrees (Finkel and Bentley, 1974), to detect and index objects in a densely populated image during the preprocessing time and tile the image based on the density of each image region (§ 4.2).
  • We propose a method to aggregate the top-$k$ rankings generated by each of the (synthetic tuple, embedding) pairs into a single final ranking while accounting for the reliability of each embedder (§ 5.1). This method enables readily adding/removing any of the current and future embedders to Needle.
  • We propose a LOF-based (Breunig et al., 2000) approach for detecting the poorly generated synthetic tuples as anomalies and discarding them (§ 5.2).
  • We conduct comprehensive experiments on multiple datasets to demonstrate the effectiveness of Needle in object detection benchmarks (§ 6.2.1), and its ability to handle scenarios with targeted queries (§ 6.2.2). We also explore the effect of different hyperparameters on Needle’s performance (§ 6.3) and the effectiveness of built-in anomaly detection method (§ 6.4).
  • Lastly, we conduct a case study where human evaluators compare Needle and OpenAI’s CLIP results, with 70% of participants preferring Needle over CLIP for their queries (§ 6.5).
2. Preliminaries
Data Model: We consider the dataset $\mathcal{D}=\{t_{1},t_{2},\ldots,t_{n}\}$ as a set of $n$ multi-modal tuples (e.g., images) with no explicit values on specific attributes or any meta-data description about them. This model complies with real-world needs, where personal devices such as cell phones have advanced cameras, enabling a vast volume of multi-modal data collection and sharing by any user of such devices.
Query Model: The input query $\varphi$ is defined as a natural language description that corresponds to a subset of tuples $\mathcal{D}_{\varphi}\subseteq\mathcal{D}$, where $\mathcal{D}_{\varphi}=\{t^{\prime}_{1},t^{\prime}_{2},\ldots,t^{\prime}_{k}\}$ represents the target subset intended for retrieval.
3. The GenAI-powered Monte-Carlo Method
Nearest-neighbor search in traditional settings is a well-studied line of research where many algorithms and indices such as Voronoi diagrams (Kolahdouzan and Shahabi, 2004), tree data structures (Yianilos, 1993), and local sensitive hashing (Jafari et al., 2021; Huang et al., 2015) have been proposed for efficient exact and approximate query answering. In such settings, the data points are specified as points in a $d$ dimensional space, i.e., $t_{i}\in\mathbb{R}^{d}$, where the goal is to find the $k$ nearest neighbors to a query point $q\in\mathbb{R}^{d}$ based on a distance function $dist:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$. That is, $k\mbox{-min}_{t_{i}\in\mathcal{D}}dist(q,t_{i})$. Such approaches, however, are not directly applicable in our setting because of the following challenges.
  1. (1)
    Multi-modal tuples such as images are not represented as points in $\mathbb{R}^{d}$.
  2. (2)
    The query is in the form of a (possibly complex) natural language phrase.
Our key idea for resolving these challenges is to leverage recent advancements in Generative AI (GenAI) and develop a randomized Monte Carlo algorithm that enables applying traditional nearest-neighbor techniques in our setting.
3.1. Vector Representations
Addressing the first challenge is possible using the vector representation (aka embedding) of the tuples (Oquab et al., 2014; Kiela and Bottou, 2014). A vector representation $\mathcal{E}:dom(t)\rightarrow\mathbb{R}^{d_{\mathcal{E}}}$ is a transformation of the tuples to a high-dimensional vector of numeric values where the semantic similarity of tuples is proportional to the cosine similarity of their vector representations. Specifically, given a vector representation $\mathcal{E}^{\ell}$, we present $\mathcal{E}^{\ell}(t_{i})$ as $\vec{v}^{\ell}_{i}=\langle v_{1},v_{2},\cdots,v_{d_{\mathcal{E}^{\ell}}}\rangle$. The distance between two vector embeddings $\vec{v}^{\ell}_{i}$ and $\vec{v}^{\ell}_{j}$ is computed based on their cosine similarity as
$\delta^{\ell}_{i,j}=\delta(\vec{v}^{\ell}_{i},\vec{v}^{\ell}_{j})=1-\cos\left(\angle(\vec{v}^{\ell}_{i},\vec{v}^{\ell}_{j})\right)$
Let $\Delta(t_{i},t_{j})$ be the (unknown) semantic distance of the tuples $t_{i}$ and $t_{j}$.
A vector representation $\mathcal{E}^{o}$ is optimal if it captures the semantic similarity of tuples, i.e., $\delta^{o}_{i,j}\sim\Delta(t_{i},t_{j}),~{}\forall t_{i},t_{j}$. Formally,
(1) $\Delta(t_{i},t_{j})\geq\Delta(t_{i},t_{k})~{}\Leftrightarrow~{}\delta^{o}_{i,j}\geq\delta^{o}_{i,k}$
In practice, however, vector representations are learned using a universe of available multi-modal data. As a result, instead of guaranteeing Equation 1, the learned embeddings satisfy a (weaker) guarantee in expectation. That is, the expected distance between the embeddings of two tuples is proportional to their semantic distance, i.e., $\mathbb{E}\left[\delta^{\ell}_{i,j}\right]\sim\Delta(t_{i},t_{j}),~{}\forall t_{i},t_{j}$.
3.2. A Baseline Approach
The state-of-the-art resolution for the second challenge is to develop vector representations that are jointly trained on the multi-modal data and textual data. Specifically, CLIP (Radford et al., 2021) is an embedding trained jointly on (images, text) pairs. As a result, CLIP transforms text and image to the vector-representation same space. Subsequently, given a natural language query, one can compute its CLIP embedding and find its nearest neighbor images in the embedding space. This approach, however, suffers from two major issues:
First, the paired text with the images are usually simple sentences (e.g., [A photo of a dog]), missing to fully describe the rich information encoded in the image. Conversely, our goal is to answer complex natural language queries, describing a scene with multiple objects and their relationships (e.g., [A dog sitting by the fireplace, staring at the cat on the couch]). As a result, as we shall verify it in our experiments (§ 6), while such embeddings may perform reasonably on simple natural language queries their performance drops significantly on complex queries.
Second, as explained in Section 3.1, learned embeddings only preserve semantic similarities in expectation. As a result, relying solely on one vector representation for query answering may be misleading and inaccurate. Therefore, we instead develop a randomized Monte Carlo algorithm that utilized a set of vector representation transformations to develop more accurate query answering.
3.3. Query Transformation using GenAI
In traditional nearest neighbor problems both the dataset tuples and the query are represented as points in the same space. However, in our problem the query is form of a natural language phrase, while the dataset contains multi-modal tuples.
Let $\varphi$ be the given natural language query, and let $\{\vec{v}^{o}_{1},\cdots,\vec{v}^{o}_{n}\}$ be the vector representations for the tuples in $\mathcal{D}$. If we could specify the corresponding vector $\vec{q}^{o}(\varphi)$ for $\varphi$ in the $\mathcal{E}^{o}$ space, we could find the $k$-nearest neighbors of $\varphi$ as
(2) $k\mbox{-NN}(\varphi,\mathcal{D})=\underset{t_{i}\in\mathcal{D}}{k\mbox{-argmin}}~{}\delta\left(\vec{q}^{o}(\varphi),\vec{v}^{o}_{i}\right)$
As a result, if we could generate the synthetic multi-modal tuple $\mathbf{g}_{\varphi}$ for which $\mathcal{E}^{o}(\mathbf{g}_{\varphi})=\vec{q}^{o}(\varphi)$, we could use Equation 2 for query answering.
Fortunately, recent advancements in generative AI, particularly the emergence of the foundation models (e.g., DALL·E 3) based on guided-diffusion techniques (Zhang et al., 2023; Rombach et al., 2022; Ramesh et al., 2022), have enabled the generation of synthetic multi-modal tuples for complex prompts $p$ (natural language descriptions of the tuple to be generated). However, due to many reasons such as the randomized nature of the foundation models, the generated tuples [^1] Throughout this paper, we refer to these synthetic tuples as “guide tuples” that represent the natural language query in search space. We use these tuples to “guide” our system toward retrieving the most relevant results. may not be the optimal synthetic tuple $\mathbf{g}_{\varphi}$, but we can assume that the generated tuple is (semantically) in the vicinity of $\mathbf{g}_{\varphi}$.
Formally, we assume every tuple generated by the foundation model for the prompt $p=toPrompt(\varphi)$ lies in an $\varepsilon$ distance from $\mathbf{g}_{\varphi}$. Where $\mathbb{E}[\varepsilon]=0$. In other words for a sample tuple $\bar{\mathbf{g}}_{\varphi}$ we have,
$\mathbb{E}\left[\Delta(\mathbf{g}_{\varphi},\bar{\mathbf{g}}_{\varphi})\right]=0$ $\Rightarrow\mathbb{E}\left[\delta\left(\mathcal{E}^{o}(\mathbf{g}_{\varphi}),\mathcal{E}^{o}(\bar{\mathbf{g}}_{\varphi})\right)\right]=0$
$\Rightarrow\mathbb{E}\left[\mathcal{E}^{o}(\bar{\mathbf{g}}_{\varphi})\right]=\mathcal{E}^{o}(\mathbf{g}_{\varphi})$
The final result follows from the principle that if the expected value of the cosine similarity between two vectors is 1, then the vectors are aligned in expectation.
3.4. The Monte Carlo Algorithm
We are now ready to develop our algorithm: a Monte Carlo randomized algorithm based on repeated sampling.
Given a natural language query $\varphi$, and at least one foundation model $\mathcal{F}$, at the first step, we generate $m$ synthetic guide tuples $\{\bar{\mathbf{g}}_{1},\cdots,\bar{\mathbf{g}}_{m}\}$. As mentioned before, we consider each guide tuple $\bar{\mathbf{g}}_{i}$ as an unbiased sample taken from a distribution with mean $\mathbf{g}_{\varphi}$.
Instead of relying on the distance estimations one vector representation, we view each embedding as a sample from the space possible transformation functions. Hence, we use multiple embedders to minimize the semantic distance estimation error. Let $\{\mathcal{E}^{1},\cdots,\mathcal{E}^{l}\}$ be the set of learned vector representations, considered for transformation. For each embedder $\mathcal{E}^{\ell}$, we assume $\mathbb{E}\left[\delta^{\ell}_{i,j}\right]=\delta^{o}_{i,j},~{}\forall t_{i},t_{j}$.
Using the generated tuples $\{\bar{\mathbf{g}}_{1},\cdots,\bar{\mathbf{g}}_{m}\}$ and the embedders $\{\mathcal{E}^{1},\cdots,\mathcal{E}^{l}\}$, we estimate the distance of each tuple $t_{i}$ to the query $\varphi$ as
(3) $\bar{\delta}_{\varphi,i}=\frac{1}{m\,l}\sum_{j=1}^{m}\sum_{\ell=1}^{l}~{}\delta\left(\mathcal{E}^{\ell}(\bar{\mathbf{g}}_{j}),\vec{v}^{\ell}_{i}\right)$
Theorem 3.1.
Let $\delta_{\varphi,i}=\delta\left(\vec{q}(\varphi),\vec{v}^{o}_{i}\right)$, $\forall t_{i}\in\mathcal{D}$. Given a positive value $\gamma$,
$\Pr\left(\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\geq(1+\gamma)\right)\vee\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\leq(1-\gamma)\right)\right)\leq\mathbf{e}^{\frac{-m\,l\,\gamma^{2}\delta_{\varphi,i}}{3}}+\mathbf{e}^{\frac{-m\,l\,\gamma^{2}\delta_{\varphi,i}}{2}}$
Proof.
Replacing $\bar{\delta}_{\varphi,i}$ with the right-hand side of Equation 3, we get
$\Pr\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\geq(1+\gamma)\right)$ $=\Pr\left(\frac{\frac{1}{m\,l}\sum_{j=1}^{m}\sum_{\ell=1}^{l}~{}\delta\left(\mathcal{E}^{\ell}(\bar{\mathbf{g}}_{j}),\vec{v}^{\ell}_{i}\right)}{\delta_{\varphi,i}}\geq(1+\gamma)\right)$
$=\Pr\left(\sum_{j=1}^{m}\sum_{\ell=1}^{l}~{}\delta\left(\mathcal{E}^{\ell}(\bar{\mathbf{g}}_{j}),\vec{v}^{\ell}_{i}\right)\geq(1+\gamma)m\,l\,\delta_{\varphi,i}\right)$
Since $\mathbb{E}\left[\delta^{\ell}_{i,j}\right]=\delta^{o}_{i,j},~{}\forall t_{i},t_{j}$ and $\mathbb{E}\left[\mathcal{E}^{o}(\bar{\mathbf{g}}_{\varphi})\right]=\mathcal{E}^{o}(\mathbf{g}_{\varphi})$,
$\mathbb{E}\left[\delta\left(\mathcal{E}^{\ell}(\bar{\mathbf{g}}_{j}),\vec{v}^{\ell}_{i}\right)\right]=\delta_{\varphi,i}$
As a result,
$\mathbb{E}\left[\sum_{j=1}^{m}\sum_{\ell=1}^{l}~{}\delta\left(\mathcal{E}^{\ell}(\bar{\mathbf{g}}_{j}),\vec{v}^{\ell}_{i}\right)\right]=m\,l\,\delta_{\varphi,i}$
Next, applying Chernoff bound, we get
$\Pr\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\geq(1+\gamma)\right)\leq\mathbf{e}^{\frac{-m\,l\,\gamma^{2}\delta_{\varphi,i}}{3}}$
Similarly,
$\Pr\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\leq(1-\gamma)\right)\leq\mathbf{e}^{\frac{-m\,l\,\gamma^{2}\delta_{\varphi,i}}{2}}$
Therefore, applying the union bound, we get
$\Pr\left(\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\geq(1+\gamma)\right)\vee\left(\frac{\bar{\delta}_{\varphi,i}}{\delta_{\varphi,i}}\leq(1-\gamma)\right)\right)\leq\mathbf{e}^{\frac{-m\,l\,\gamma^{2}\delta_{\varphi,i}}{3}}+\mathbf{e}^{\frac{-m\,l\,\gamma^{2}\delta_{\varphi,i}}{2}}$
Using the estimated distances, we identify the $k$-nearest neighbors of $\varphi$ as
$\overline{k\mbox{-NN}}(\varphi,\mathcal{D})=\underset{t_{i}\in\mathcal{D}}{k\mbox{-argmin}}~{}\bar{\delta}_{\varphi,i}$
4. System Development
Figure 2. Needle Architecture
Having discussed the theoretical foundation of our system, we can now explain its architecture and different components in this section. While our GenAI-powered Monte Carlo method is modality agnostic, we will explain the development of our system for Image Retrieval. The overall architecture of Needle is illustrated in Figure 2. The architecture can be divided into two main parts: components involved in Preprocessing [§ 4.1, § 4.2, § 4.3] and Inference [§ 4.4, § 4.5, § 4.6, § 4.7].
4.1. Preprocessing
The preprocessing phase involves a one-time operation to create vector representations, also known as embeddings, of the input dataset using a set of predefined embedders. Since images are typically regarded as being strongly associated with their representing objects(Zhao et al., 2019), embedding models may underperform when applied to images containing numerous small objects (Liu et al., 2021b).
To address this limitation, we employ a tiling strategy that divides the input image into multiple tiles based on the density of objects in each region (§ 4.2).
Each of these tiles is then processed by the embedders, and the resulting embeddings are stored on disk. This process can be executed in parallel for each embedder, with performance significantly enhanced by using multiple GPUs to handle tile embedding.
4.2. Smart Tiling
Figure 3. An Illustration of Smart Tiling on an image with lots of objects
Algorithm 1 Smart Tiling
1:image, d, objects_borders, min_size
2:image tiles
3:$(h,w)\leftarrow\text{size of }image$
4:tiles [$\leftarrow$[]← [ ]
5:function CountObjects(x, y, w, h)
6:    return number of unique objects in the region $(x,y,w,h)$
7:end function
8:
9:function Split(x, y, w, h)
10:    num_objects $\leftarrow$ CountObjects(x, y, w, h)
[…]
12:         Append((x, y, w, h)) to tiles
13:    else
14:         Compute half_w, half_h
15:         Calculate objs_left, objs_right, objs_top, objs_bottom
16:         if vertical split is preferred & aspect ratio is reasonable then
17:             Split(x, y, half_w, h) $\triangleright$ Split vertically
18:             Split(x + half_w, y, half_w, h)
19:         else
20:             Split(x, y, w, half_h) $\triangleright$ Split horizontally
21:             Split(x, y + half_h, w, half_h)
22:         end if
23:    end if
24:end function
25:
26:Split(0, 0, w, h)
27:return tiles
To capture all the details within an image, it is important to focus on densely populated regions. Therefore, we should split each image into different tiles based on the density. In addition, most embedders are trained on square images (He et al., 2016; Simonyan and Zisserman, 2014; Szegedy et al., 2017) and perform optimally when applied to such shapes (Shahnaz and Mollah, 2022; Azulay and Weiss, 2019), which means excessively tall or narrow tiles should be avoided. In order to achieve these two goals, inspired by (Finkel and Bentley, 1974), we propose our Smart Tiling procedure. We employ an edge detection technique (Canny Edge Detection (Canny, 1986)) to identify object boundaries within the image and recursively split the image to balance the distribution of objects across tiles as effectively as possible. The objective of this algorithm is to divide the image into smaller tiles, where each tile contains no more than a specified number of objects $d$, while maintaining a balanced aspect ratio.
The algorithm recursively splits the image by first checking if a region contains more objects than the allowed threshold $d$. If this condition is met, the region is split either vertically or horizontally. The algorithm calculates the object distribution for potential splits and selects the one that provides the most balanced distribution of objects across the two resulting tiles. To avoid generating disproportionate tiles, the algorithm also checks the aspect ratio and favors splits that maintain near-square dimensions. This recursive process continues until each tile satisfies the object limit constraint or it reaches the minimum acceptable width/height, at which point the tiles are finalized. Algorithm 1 presents the pseudocode for our Smart Tiling approach.
4.3. Vector Store
Once the raw image data is processed into vectors, it is indexed within the vector store. The vector store provides a maximum inner product search interface with low latency, which is crucial since users expect prompt results from the system. In our implementation, we employ hnswlib, which offers approximate maximum inner product search capabilities. This is consistent with the functionality provided by most vector stores.
4.4. Query Refinement
Having discussed the components involved in Preprocessing, we now explore the components used during Inference. After receiving the query in natural language text, Needle refines the input query into a suitable prompt. The first step is to remove punctuation and redundant characters from the query. Additionally, Needle supports more advanced query refinement techniques.
[…]
In addition, given that current Foundation Models have limitations in fully understanding some types of prompts (Liu et al., 2022; Feng et al., 2022; Lee et al., 2023; Liu et al., 2024), our system can employ various query refinement techniques (Liu et al., 2021a; Liu and Chilton, 2022; Wang et al., 2023; Witteveen and Andrews, 2022) to paraphrase and reformat the input query. These transformations help better align the generated query tuples with the original intent. These techniques help overcome prompt misinterpretations and improve the relevance of generated results.
4.5. Foundation Model and Guide Image Generation
After refining the query, We need to generate guide tuples representing the query using foundation models. Needle can use multiple (black-box) foundation models to generate guide images, with the number of guide images fine-tuned as a hyperparameter, denoted as $m$. Foundation models are capable of producing images with varying levels of quality and resolution. However, for the purposes of Needle, high-quality realistic images are unnecessary to achieve accurate search results. Lower-resolution images are sufficient, provided that they closely resemble the input query. In our experiments (§ 6.3), we will explore the impact of $m$, the size of the generated images, and the overall quality of the foundation model in the final results.
4.6. User Interface and Feedback Process
Figure 4. Needle User Interface
The current implementation of Needle offers a web-based user interface that allows users to search based on their queries (Figure 4). Users can interact with the system in two modes: Feedback mode (on) or No-Feedback mode (off). When Feedback mode is enabled, the guide tuples are presented to the user prior to the search process. Users can evaluate the quality and relevance of the generated guide tuples with respect to their query, which can assist Needle in retrieving more relevant images. In No-Feedback mode, Needle automatically identifies and discards outlier tuples from the generated images without user intervention (§ 5.2).
4.7. Embedders and Rankings
After generating the guide images, Needle utilizes predefined embedders to convert the guide tuples into vector representations and performs a k-NN lookup on the vector stores (§ 4.3). The score of each image is determined by the maximum score among all of its tiles (with the entire image also treated as a tile). Subsequently, the results are aggregated using our aggregation mechanism (§ 5.1) to produce a final ranking, which is then presented to the user.
5. Practical Optimization
In this section, we discuss our practical optimizations to improve the efficiency of our system.
First, we address the issue of aggregating results from multiple embedders into a unified ranking (§ 5.1), followed by a discussion on detecting anomalies in generated synthetic guide tuples (§ 5.2).
5.1. Rankings Aggregation and Embedder Trust Mechanism
Needle has been designed with scalability and flexibility in mind, ensuring it is not restricted to any specific embedder. Particularly considering the rapid advancements in vector representations, our system is designed to dynamically add, remove, or upgrade embedders. This will allow our system to enhance its performance as more advanced embedders become available.
Needle aggregates the results of the available embedders in real-time to achieve optimal outcomes. Furthermore, Needle monitors the performance of embedders for different queries, adjusting its trust in them based on their historical performance. The performance of embedders may vary depending on the type of image input. For instance, one embedder may excel at distinguishing different breeds of dogs but underperform in fruit classification tasks, while another may perform better in the latter. Needle addresses this variability by dynamically adjusting the reliability score of embedders for various query topics.
Consider a (predefined) set of topics $T$, and let the natural language query $\varphi$ belong to the topic $t=t(\varphi)\in T$ [^2] The topic $t$ of query $\varphi$ is identified using a topic classification mechanism or is part of the query input.. Let $\mathbf{G}=\{\bar{\mathbf{g}}_{1},\cdots,\bar{\mathbf{g}}_{m}\}$ be the set of guide images generated for $\varphi$ and consider the embedders $\mathbf{E}=\{\mathcal{E}^{1},\cdots,\mathcal{E}^{l}\}$. For a guide image $\bar{\mathbf{g}}_{j}$, let $R^{i}_{j}=\mathcal{R}^{i}(\varphi,\bar{\mathbf{g}}_{j})$ be the sorted list of $k$ nearest images in $\mathcal{D}$ to the guide image $\bar{\mathbf{g}}_{j}$ according to the embedder $\mathcal{E}^{i}$.
Let $R^{i}_{j}$[r] be the $r$-th element of $R^{i}_{j}$.
For each topic $t\in T$, we assign a weight $w_{i}^{t}\in[0,1]$ to each embedder $\mathcal{E}^{i}$, representing the reliability of the embedder for answering the queries within topic $t$. The weights are normalized so that for each topic $t$,
$\sum_{i=1}^{l}w_{i}^{t}=1.$
5.1.1. Aggregation
Given the ranked lists from all the embedders and the topic $t$ of the query $\varphi$, the objective is to aggregate these lists into a single final ranking $R_{\varphi}$ by incorporating the topic-specific reliability weights $w^{t}_{i}$.
Let $I_{h}\in\mathcal{D}$ be a tuple returned by at least one of the embedders. Let $rank(I_{h},R^{i}_{j})$ be the rank assigned to $I_{h}$ according to the list $R^{i}_{j}$. Let $\mathsf{S}:[k]\rightarrow[0,1]$ be a monotonically decreasing function that specifies the importance of each rank position in the top-$k$. In particular, while being agnostic to the choice of $S(.)$, following the literature (Cummins, 2012; Arampatzis and Robertson, 2011; Faria et al., 2010), we adopt the following position-importance function in Needle:
$\mathsf{S}(i)=\begin{cases}\frac{1}{i}&~{}i\leq k\\ 0&~{}\text{otherwise}\end{cases}$
Then, the aggregated score of $I_{h}$ is computed as:
(4) $score(I_{h})=\sum_{i:\mathcal{E}^{i}\in\mathbf{E}}w^{t}_{i}\cdot\mathsf{S}(rank(I_{h},R^{i}_{j}))$
The top-$k$ tuples based on Equation 4 are returned as the query output, denoted as $R_{\varphi}$.
5.1.2. Dynamic Reliability-weight Adjustment
Suppose after showing the top-$k$ results $R_{\varphi}$ to the user, we receive feedback from him/her, which is a subset of results marked as irrelevant to the query $\varphi$  [^3] We acknowledge that such feedback may not commonly be available after the system deployment. However, crowdsourced feedback on a small set of topic-based benchmarks can be employed to calculate these weights before deploying a new embedder to our system..
Let $\mathcal{T}=\{I^{\prime}_{1},I^{\prime}_{2},\dots,I^{\prime}_{s}\}$ be the set of irrelevant tuples where $\mathcal{T}\subseteq R_{\varphi}$
Using this feedback, we update the reliability weights of the embedders for each topic using the Multiplicative Weight Update Method (MWUM) (Arora et al., 2012). Let $t=t(\varphi)$ be the topic of the query $\varphi$. Then the partial loss of the embedder $\mathcal{E}^{i}$ for a topic $t$ based on $\varphi$ is computed as,
(5) $\Delta loss_{\varphi}(t,i)=\sum_{j:\bar{\mathbf{g}}_{j}\in\mathbf{G}}\sum_{I^{\prime}_{h}\in\mathcal{T}}\mathsf{S}(rank(I^{\prime}_{h},R^{i}_{j}))$
We only penalize those embedders that highly ranked the irrelevant results within the query topic $t$.
The topic-specific weights are updated multiplicatively:
$w_{i}^{t}(t+1)=w_{i}^{t}(t)\cdot(1-\eta\cdot\Delta loss_{\varphi}(t,i))$
where $\eta\in(0,1)$ is the learning rate controlling how much the trust in each embedder changes over time. Finally, the weights are normalized for topic $t$ to ensure $\sum_{i=1}^{l}w_{i}^{t}=1$.
5.2. Anomaly Detection in Query Images
In addition to rankings aggregation, Needle employs an outlier detection mechanism to identify potentially anomalous guide images based on their embeddings from multiple embedders. These anomalies might indicate data quality issues or images that are inherently different from the majority. This process combines the embeddings from all available embedders and uses a weighted approach to adjust for the reliability of each embedder.
We are given a set of $m$ guide images $\mathbf{G}=\{\bar{\mathbf{g}}_{1},\bar{\mathbf{g}}_{2},...,\bar{\mathbf{g}}_{m}\}$ and a set of $l$ embedders $\mathbf{E}=\{\mathcal{E}^{1},\mathcal{E}^{2},...,\mathcal{E}^{l}\}$. For each image $\bar{\mathbf{g}}_{j}$, the embedder $\mathcal{E}^{i}$ produces a high-dimensional embedding $\vec{v}^{i}_{j}=\mathcal{E}^{i}(\bar{\mathbf{g}}_{j})$. Thus, for each image $\bar{\mathbf{g}}_{j}$, we have $l$ corresponding embeddings from the embedders:
$\{\vec{v}^{1}_{j},\vec{v}^{2}_{j},\dots,\vec{v}^{l}_{j}\}.$
Our goal is to detect if any guide image $\bar{\mathbf{g}}_{j}$ is an anomaly by analyzing the embeddings across all embedders.
5.2.1. Dimensionality Reduction using UMAP (McInnes et al., 2018)
Given the high-dimensional nature of the embeddings, we first apply dimensionality reduction to project the embeddings into a lower-dimensional space. For each embedder $\mathcal{E}^{i}$, we reduce the dimensionality of the embeddings $\{\vec{v}^{i}_{1},\vec{v}^{i}_{2},\dots,\vec{v}^{i}_{m}\}$ to a lower dimension $d$ using UMAP (Uniform Manifold Approximation and Projection) (McInnes et al., 2018).
UMAP constructs a weighted $k$-nearest neighbors graph of the high-dimensional points, optimizes the layout in a low-dimensional space, and minimizes the cross-entropy between the two spaces. The reduced embeddings are given by:
$\vec{v^{\prime}}^{i}_{j}=\text{UMAP}(\vec{v}^{i}_{j})\in\mathbb{R}^{d}$
where $d\ll\text{dim}(\vec{v}^{i}_{j})$.
5.2.2. Outlier Detection using LOF (Breunig et al., 2000)
After dimensionality reduction, we apply the Local Outlier Factor (LOF)(Breunig et al., 2000) algorithm to identify outliers. LOF assigns an outlier score to each data point based on the density of its neighborhood. The LOF score for image $\bar{\mathbf{g}}_{j}$ with respect to embedder $\mathcal{E}^{i}$ is defined as:
$\text{LOF}^{i}(\bar{\mathbf{g}}_{j})=\text{LOF}(\vec{v^{\prime}}^{i}_{j})$
This score indicates how much of an outlier $\bar{\mathbf{g}}_{j}$ is in the context of the embeddings from embedder $\mathcal{E}^{i}$, with higher scores representing a greater likelihood of being an outlier.
5.2.3. Aggregation of Outlier Scores
The final outlier score of an guide image $\bar{\mathbf{g}}_{j}$ is computed by aggregating the individual scores, weighted with the reliability weight of each embedder:
$S(\bar{\mathbf{g}}_{j})=\sum_{i=1}^{l}w_{i}\cdot\text{LOF}^{i}(\bar{\mathbf{g}}_{j})$
5.2.4. Anomaly Detection Threshold
In the end, if the outlier score $S(\bar{\mathbf{g}}_{j})$ of an image $\bar{\mathbf{g}}_{j}$ exceeds a given threshold $\tau$, the image is flagged as an anomaly:
$S(\bar{\mathbf{g}}_{j})>\tau\quad\Rightarrow\quad\bar{\mathbf{g}}_{j}\text{ isan anomaly}$
The threshold $\tau$ is a hyper-parameter adjusted according to the sensitivity required by the user. Higher values of $\tau$ promote greater diversity in the retrieved images, potentially increasing recall but risking the inclusion of images that do not adequately represent the query. This may result in lower accuracy and decreased precision. In contrast, lower values of $\tau$ force Needle to adhere more closely to the query, leading to higher precision in the results but potentially missing some relevant images.
6. Experiments
This section presents results of our experiments designed to evaluate the effectiveness of our proposed system, Needle. We begin by outlining the experimental setup (§ 6.1) and proceed with a proof of concept (PoC) (§ 6.2), where we demonstrate Needle ’s ability to retrieve relevant images across different scenarios.
In the first scenario, we focus on object detection tasks, which may involve identifying either rare (single occurrence) or common objects within the dataset (§ 6.2.1). In the second scenario, the goal is to locate a specific image using its natural language description. We evaluate how accurately Needle can retrieve the target tuple based on the provided description.
Following the PoC, we explore the impact of various hyper-parameters on Needle’s performance (§ 6.3) and how can Needle perform well, even with a few low-resolution guide tuples. We then analyze the performance of the anomaly detection component embedded within Needle and will compare it with human evaluation results(§ 6.4). Finally, we present a case study in which human evaluators compare Needle ’s results with CLIP (§ 6.5).
6.1. Experimental Setup
6.1.1. Datasets
We have used Caltech256 (Griffin et al., 2022), LVIS (Gupta et al., 2019), and the Berkeley DeepDrive (BDD) (Yu et al., 2020) datasets for object detection. Caltech256 is a relatively simple benchmark that includes 256 categories of objects with 30,000 images. LVIS is a long-tailed benchmark with over 1,300 categories and 100,000 images. Due to the fact that some categories in LVIS have fewer than 10 images across the entire dataset, this benchmark is considered particularly challenging for object detection tasks. BDD is a large-scale, diverse dataset for autonomous driving, which contains over 100,000 video frames and images annotated for object detection, lane detection, and drivable area segmentation.
For the targeted queries experiments, we use two image captioning benchmarks: Nocaps (Agrawal et al., 2019) and SentiCap (Mathews et al., 2016). Nocaps is a large-scale image captioning dataset that extends the Open Images dataset (Kuznetsova et al., 2020), featuring captions for novel object categories and their relations to each other. SentiCap (Mathews et al., 2016), on the other hand, focuses on sentiment-driven image captions, containing images annotated with captions that convey either positive or negative sentiments, making it suitable for evaluating models on emotional and subjective image descriptions.
6.1.2. Evaluation Metric
In our experiments, we report Average Precision (AP) as the primary evaluation metric. AP is a widely used accuracy metric in information retrieval, as it places greater emphasis on retrieving relevant results earlier in the ranking compared to simpler metrics such as precision or F1-score. AP values range from 0 to 1, with an AP of 1 indicating perfect accuracy, meaning all relevant results appear before any irrelevant ones. For a given query, AP is calculated as the average of the individual precision scores at each possible recall threshold: $AP=(\sum_{i=1}^{R}P_{i})/R$, where $R$ is the total number of relevant results in the dataset, and $P_{i}$ represents the precision when considering the top $i$ relevant results. In addition to reporting the mean Average Precision (mAP) as the final performance indicator for each baseline, we also report Hit Rate for targeted queries. Since these queries are designed to find exactly one target image, Hit Rate measures whether the relevant image appears in the results, regardless of its rank. The mean Hit Rate (mHR) across all queries is also presented for each system.
6.1.3. Baselines
In our experiments, we use the latest and highest-performing version of CLIP, specifically $ViT$-$B/32$, as the baseline model (Radford et al., 2021). CLIP (Contrastive Language–Image Pretraining) is a powerful vision-language model developed by OpenAI, designed to learn visual concepts from natural language descriptions. It achieves this by jointly training an image encoder and a text encoder using a contrastive learning objective, aligning images and their corresponding textual descriptions in a shared embedding space. CLIP has demonstrated strong zero-shot performance across a wide range of visual tasks, making it a robust baseline for image retrieval tasks.
In all experiments, we will report AP for all queries in the dataset and their “hard subset”. For each dataset, we define the hard subset as the set of tuples for which CLIP AP is less than 0.5.
6.1.4. Environmental Setup
All experiments were conducted on a local server equipped with 128 GB of memory, 32 CPU cores, and two NVIDIA GeForce RTX 2080 Ti GPUs. Four different Foundation Models were utilized throughout the experiments. runwayml-v1.5 and sdxl-turbo-v1.0 were hosted locally on the aforementioned server, while DALL·E 3 and Flux-Schnell were accessed via paid APIs.
6.2. Proof of Concepts
6.2.1. Object Detection Experiments
Table 1. Performance (mAP) Comparison Between CLIP and Needle on Object Detection Benchmark Datasets
Dataset (#hard/#all) All Queries Hard Subset
CLIP Needle CLIP Needle
BDD (2/11) 0.76 0.83 0.01 0.24
Caltech256 (10/256) 0.94 0.95 0.26 0.59
LVIS (1049/1203) 0.16 0.31 0.08 0.23
Average 0.62 0.70 0.12 0.35
((a)) LVIS
((b)) Caltech256
((c)) BDD
Figure 5. Distribution of Average Precision for different queries for CLIP and Needle, split into hard and easy subsets.
In these experiments, we demonstrate the effectiveness of Needle in basic object detection tasks. We utilize state-of-the-art benchmark datasets, LVIS (Gupta et al., 2019), Caltech256 (Griffin et al., 2022), and BDD (Yu et al., 2020), for this evaluation. For each query, the number of retrieved images is set to 60 ($k=60$), and we aim to identify 10 relevant tuples within the results ($R=10$). Two local Stable Diffusion models are employed as Foundation Models: runwayml-v1.5 and sdxl-turbo-v1.0 with guide image resolution of $768\times 768$. For each model, two guide images are generated, resulting in a total of four guide images per query ($m=4$). Table 1 presents a comparative analysis of performance between CLIP and Needle across the different datasets. Notably, Needle consistently demonstrates superior performance across all categories, in particular, for the hard subsets, Needle achieves around 200% improvement in AP compared to CLIP on average, highlighting its effectiveness in addressing more challenging object detection scenarios. Figure 5 shows the AP distribution of CLIP and Needle for different datasets. Needle demonstrates significant improvements in AP, even for queries where CLIP initially performs very poorly, with AP close to zero.
It is important to note that all generated guide images are produced using local Stable Diffusion models, which generate lower quality images than commercial alternatives. This result highlights that Needle imposes much looser constraints on the quality of guide tuples than humans do. Even when the guide images are distorted or lack realism, Needle can robustly retrieve relevant images using these guide tuples.
6.2.2. Targeted Queries
Figure 6. Illustration of different specific queries, the generated guide images and the comparison of CLIP and Needle results
Table 2. Performance (mAP, mHR) Comparison Between CLIP and Needle on Targeted Queries Benchmark Datasets
Dataset (#queries) mAP mHR
CLIP Needle CLIP Needle
Nocaps (100) 0.11 0.37 0.55 0.73
SentiCap (100) 0.24 0.34 0.71 0.77
Average 0.18 0.36 0.63 0.75
Needle excels at handling targeted queries and identifying rare tuples. We demonstrate this by running two image captioning benchmarks, Nocaps (Agrawal et al., 2019), and SentiCap (Mathews et al., 2016), using both CLIP and Needle. For each experiment, the systems are provided with a query that describes a specific image in the dataset, with the search engines returning 10 images. For this task, we enhance Needle with two more powerful and precise Foundation Models: DALL·E 3  [^4] https://openai.com/index/dall-e-3/ and Flux-Schnell  [^5] https://replicate.com/black-forest-labs/flux-schnell. These models are selected due to the need for guide images that accurately represent the detailed input queries, as these queries include significant complexity and detail.
In addition to reporting mean Average Precision (mAP), we also evaluate the mean Hit Rate (mHR). Hit Rate is particularly useful in cases where the number of retrieved images is small, and the focus is on finding the correct image rather than its specific ranking.
As shown in Table 2, Needle significantly outperforms CLIP in locating the target image. Needle successfully retrieves the correct image in over 75% of queries within the top 10 results, marking around 20% improvement over CLIP’s performance in mHR and 100% improvement in mAP. Figure 6 presents four different queries from the Nocaps (Agrawal et al., 2019) dataset, along with the generated guide images for Needle and the top-1 results from both CLIP and Needle. As illustrated, CLIP frequently overlooks key aspects of the input prompt and focuses on a single keyword rather than capturing the full intent of the query. In contrast, Needle effectively captures all inherent complexity of query within its guide tuples and identifies the target image using only two guide images.
6.3. Parameter Tuning
((a)) Runwayml (Caltech256)
((b)) sdxl-turbo (Caltech256)
((c)) Runwayml (LVIS)
((d)) sdxl-turbo (LVIS)
Figure 7. Heatmaps of Needle mAP for two different datasets (Caltech256 and LVIS) using two local Foundation Models (Runwayml and sdxl-turbo) while altering number of Generated Images $m$ and Generated Image Size
In this section, we investigate the effect of varying the number of generated guide images ($m=\{1,2,4\}$), the image resolution ($\{512\times 512,768\times 768,1024\times 1024\}$), and the performance of different foundation models (runwayml, sdxl-turbo) on Needle’s effectiveness. The experiments are conducted on both the Caltech256 and LVIS datasets. Figure 7 illustrates the mAP values for different combinations of these parameters. As a general trend, increasing the number of images and the resolution improves mAP performance. However, there are deeper insights to consider: first, sdxl-turbo exhibits more consistent and slightly better performance across different image sizes compared to runwayml; second, while increasing the number of generated images $m$ improves results, the improvement diminishes and stabilizes after a certain point. A similar trend is observed for image sizes. These results suggest that Needle can maintain high performance even with a limited number of generated images and lower resolution inputs.
6.4. Feedback Mode
((a)) LVIS
((b)) BDD
Figure 8. Comparison of AP distribution for Needle on Feedback mode on and off
CLIPNeedleBothNeither002020202040404040606060608080808010010010010023.2323.2323.2323.2352.5252.5252.5252.5214.1514.1514.1514.1510.110.110.110.1Percentage (%)
Figure 9. Case study results
Table 3. Comparison of Needle’s performance (mAP) for different datasets with Feedback mode enabled (ON) and built-in Anomaly detection (OFF)
Dataset (#hard/#all) All Queries Hard Subset
OFF ON OFF ON
BDD (2/11) 0.82 0.84 0.22 0.25
LVIS (84/100) 0.28 0.30 0.19 0.21
Average 0.55 0.57 0.21 0.23
In this subsection, we investigate the impact of enabling the feedback procedure in Needle on mAP. A human evaluator was employed to verify the quality and alignment of the generated images with the input query prior to retrieving results. For evaluation, we used LVIS and BDD datasets, generating four tuples per query ($m=4$) with an image size of $768\times 768$, utilizing the runwayml foundation model. We compare the mAP of Needle’s built-in anomaly detection method (§ 5.2) to the performance when human feedback is incorporated.
Figure 8 presents the AP distribution for both baselines. As shown, there is a slight improvement when feedback mode is enabled, primarily raising queries with $AP\approx 0.4$ in LVIS and $AP\approx 0.8$ in BDD toward higher APs. This indicates that: (a) Needle’s built-in anomaly detection method performs effectively in practice, achieving results comparable to a human evaluator, and (b) human intervention is most beneficial for queries that already have a moderate AP, whereas queries for which foundation models fail to generate suitable images do not significantly improve with feedback alone.
6.5. Case Study
To further validate Needle’s practical applicability, we conducted a case study with human evaluators. We utilized the LVIS dataset as our primary image repository, as it contains over 100k images spanning various categories, ensuring that any randomized query would likely have related images. This made LVIS an ideal choice for our case study.
Each human evaluator was asked to write five queries of their choice, with varying levels of complexity. For each query, we retrieved 20 relevant images from both CLIP and Needle, displaying the results side by side with randomized orientations for each query. Evaluators were then asked to select the more relevant results, choosing from four options: ”Left is better,” ”Right is better,” ”Both are good enough,” and ”Neither is good enough.”, then we recorded the engine behind that option as preferred engine.
A total of 20 human evaluators participated in the study, contributing 100 queries. The queries ranged from simple ones like ”Celery stew” to more complex descriptions such as ”A person walking on the sidewalk of a river in Fall at Chicago.”, Figure 9 presents the results of this user study. If we consider only the queries where one engine is preferred over the other, Needle is chosen as the better engine in almost 70% of cases.
7. Related Work
Our paper relates to work in the following areas:
Multi-modal Data Retrieval: Multi-modal data retrieval, which integrates various data types (e.g., text, images, video, and audio), has attracted significant attention. Early methods focused on shared representations by mapping modalities into a common latent space using canonical correlation analysis (CCA) (Rasiwasia et al., 2010; Wang et al., 2015b) and kernel-based techniques (Jia et al., 2019). Recent approaches leverage deep learning models, including cross-modal encoders (Wang et al., 2015a; Liu et al., 2017) and joint embedding networks (Balaneshin-kordan and Kotov, 2018; Rubio et al., 2017). Pre-trained transformers like CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) highlight the effectiveness of combining vision and language models for cross-modal tasks.
Advanced techniques, such as attention mechanisms (Yang et al., 2021; Tang et al., 2023) and contrastive learning (Pham et al., 2024), further enhance retrieval by capturing fine-grained relationships. Feedback mechanisms (Moll et al., 2023; Zhou and Huang, 2003) and user preferences (Zhang and Zhang, 2010; Chen et al., 2020) improve multi-modal search results. Research also extends to other modalities, such as text-video retrieval (Li et al., 2019; Markatopoulou et al., 2017; Torabi et al., 2016; Mithun et al., 2018; Miech et al., 2019; Luo et al., 2021) and text-audio retrieval (Xin et al., 2023; Chechik et al., 2008; Koepke et al., 2022). For a detailed survey of cross-modal retrieval methods and future directions, see (Li et al., 2023).
GenAI in Data Management: The intersection of Generative AI (GenAI) and data management has emerged as a promising area, driving innovations in how data is generated, managed, and utilized. GenAI models, such as GPT (Achiam et al., 2023) and Stable Diffusion (Rombach et al., 2022), have been employed for synthetic data generation to augment datasets (Tanaka and Aranha, 2019; Shahbazi et al., 2024), improve model training (Le et al., 2017; Tremblay et al., 2018), and address data scarcity or imbalance issues (Erfanian et al., 2024). In data cleaning and integration, these models assist with tasks like imputation (Villaizán-Vallelado et al., 2024), and anomaly detection by generating plausible data patterns (Ye et al., 2024). Recent works also explore the role of GenAI in automating query generation (Trummer, 2022, 2023; Li and Xie, 2024) and data summarization (Dar et al., 2024), enhancing the efficiency of database systems and analytics platforms.
Natural Language Query Processing: The evolution of natural language to structured query translation has advanced from rule-based systems to sophisticated large language models (LLMs). Early methods used template matching and predefined rules (Androutsopoulos et al., 1995; Li and Jagadish, 2014), which were inflexible and required frequent manual updates. Machine learning approaches introduced more adaptability by learning query patterns from annotated datasets (Ma et al., 2020; Zhang et al., 2019), though they struggled with complex language constructs.
Deep learning models, especially transformers like BERT (Devlin, 2018) and GPT (Brown, 2020), enabled semantic understanding and laid the foundation for text-to-SQL systems (Li et al., 2024). Modern LLMs, such as PaLM (Chowdhery et al., 2023) and GPT-4 (Achiam et al., 2023), generate context-aware queries across multiple domains with minimal supervision. Retrieval-augmented generation (RAG) further enhances query precision by leveraging external knowledge (Lewis et al., 2020). Ongoing efforts focus on fine-tuning LLMs to improve accuracy and efficiency in converting natural language to structured queries (Trummer, 2022, 2023; Li and Xie, 2024).
8. Discussions and Limitations
8.1. Reliance on Existing Models
Needle extensively relies on Foundation Models for guide tuple generation and on embedders for generating vector representations of images. Both these components have their own capabilities, limitations, and inherent biases. This reliance presents both strengths and weaknesses in Needle’s architecture. The strength lies in the ability to easily improve performance by upgrading the internal Foundation Models and embedders. However, the limitations of these models also become limitations for Needle. To mitigate this, we designed Needle to support multiple Foundation Models and embedders, enabling it to draw from a broader range of knowledge and reduce the impact of individual model limitations.
8.2. Efficiency
We assess the efficiency of Needle across two phases: Preprocessing and Inference.
Preprocessing: In the preprocessing phase, two highly parallelizable tasks are performed: tiling and embedding creation. Tiling is a fast operation, but it can linearly increase the time required for embedding creation. To improve time efficiency, one can increase the minimum tile width and height, thereby reducing the number of tiles. The time required for embedding creation can be significantly improved by utilizing multiple GPUs. Since the preprocessing phase is a one-time operation that can be easily parallelized and fine-tuned, it does not present a major concern for Needle’s overall performance.
Inference: Needle’s inference consists of three main sequential operations: (a) generating guide tuples, (b) k-NN lookup, and (c) ranking aggregation. For an input dataset of size $n$, retrieving $k(1\leq k\leq n)$ results with constants $m$ (number of guide images) and $l$ (number of embedders), (a) is a constant-time operation with respect to $n$ and $k$. Step (b)’s time complexity depends on the underlying vector store algorithm, and for HNSW data structure (Malkov and Yashunin, 2018), is $O(k\log(n))$. Step (c) has a time complexity of $O(k)$ using Fagin’s Algorithm (Fagin et al., 2001).
In our experiments, where the number of tuples (tiles) is at most 400,000 for any k, (b) k-NN lookup and (c) the rankings aggregation take less than one second using current approximate vector stores. The bottleneck of the entire process lies in (a), generating guide tuples. The good news is that this operation is independent of the dataset size $n$ and number of images to retrieve $k$, and only relies on the number of guide images $m$. Moreover, the generation process can run in parallel, significantly improving performance.
To summarize, for small values of $n$, Needle’s inference time may be slower than other methods due to the constant time required for guide image generation. However, as $n$ increases, the effect of step (a) diminishes, and Needle’s efficiency becomes comparable to other methods.
8.3. Extensions to Other Modalities
The core idea behind Needle is adaptable to various modalities beyond images, including audio, video, and others. However, due to the current limitations in publicly available foundation models for audio and video generation, our focus in this paper has been restricted to the image modality. As foundation models for these additional modalities continue to improve and become more accessible, we plan to extend Needle to support them. This will allow Needle to operate across multiple data types, unlocking further potential for handling complex, multi-modal queries in the near future.
9. Conclusion
In this paper, we took a step towards addressing the challenge of answering complex natural language queries in multi-modal data settings by proposing a Generative-AI Powered Monte Carlo method. Our proposed method utilizes foundation models to generate synthetic guide tuples that capture the complexities of the input query; then, we leverage these guide tuples to represent the natural language query in multi-modal data space and apply traditional nearest neighbor methods.
Following this method, We implemented Needle for image retrieval. We proposed different solutions for practical challenges in our system’s development. Our comprehensive experiments demonstrated that our system, Needle, significantly outperforms the state-of-the-art techniques in image retrieval.
While our proposed technique is not limited to image datasets, its development for other modalities, such as audio and video, requires foundation models that are currently not publicly available. Upon the availability of such models, we aim to extend the scope of Needle to a wider range of modalities in our future work.